A telephone call center where people place marketing calls to customers has a probability of success of 0.06. Use the binomiaWhat is the probability that placing 50 marketing calls would result in generating at least 4 new customers? 0.8100 0.4162 0.0.1900 0.1794What is the probability that placing 50 marketing calls would result in generating 2 or less new customers? 0.6473 0.3527 0.80.8206 0.1794 0.5838 Question 11 (7 points) What is the minimum number of marketing calls that need to be placed in order to150 130 170 50
Answer #1

Question re ; X=0,1,2, - 50 þ=0.06 ;q=1-b=1-0106 = 0.94 ;*=50 p(x)= Ge pr. qnx ; x =0,), 3 - - ., p(x)= 50G (0.06) (6.94) 5Pr (X<2) = Px (x = 0) + f(x=1)+ Pe (X=2) 5066.06)%.94)50+50660.06) (0.94) 94 506,66-06) (0-34) 18 50 49 - (0.94) 3°+(50) (0